OpenStudy (anonymous):

A force of 526 lb is required to pull a boat up a ramp inclined at 23˚ with the horizontal. How much does the boat weight?

OpenStudy (p0sitr0n):

acceleration in the ramp = g*sin (x) = g * sin (23) means m*g*sin(26)=526 m = 526 / (g) (sin26) calculate it